Determine the number of ways to form 4-digit numbers by using the digits 1,2,3,4,5,6, and 7 if repetitions of digits are allowed.
Answer by stanbon(75887)

Determine the number of ways to form 4-digit numbers by using the digits 1,2,3,4,5,6, and 7 if repetitions of digits are allowed.
---
Ans: 7*7*7*7 = 7^4 = 2401 numbers
============
Cheers,
Stan H.